OpenStudy (ashley1nonly):

In exercises the component form of the vector. 12.)The sum of AB and CD where A(1,-1) B(2,0) C(-1,3) D(-2,2) I got <0,4> Is this the correct answer?

OpenStudy (ashley1nonly):

AB=<3,-1> CD=<-3,5> So adding i got <0,4>

OpenStudy (jdoe0001):

hmmm does it mean \(\large A\cdot B+C\cdot D?\)

OpenStudy (ashley1nonly):

They are vectors

OpenStudy (jdoe0001):

\(A\cdot B\ne A+B\) thus

OpenStudy (jdoe0001):

I could be wrong, but sounds as if it's asking for the "dot product" of AB and the "dot product" of CD and then their sum

OpenStudy (ashley1nonly):

No that's in the next section. This is just asking for the sum of the two vectors giving these two points and i wanted to know if it was right.

OpenStudy (jdoe0001):

hmmm